# Analyzing the dense matter equation of states in the light of the compact object HESS J1731-347

### v1

@@ III.2 Strange EoS
Figure 5: Odds ratio plot of the strange matter EoS with $x_{v}=0.6$ and B = 139.5 with other strange matter EoSs. The color scheme is the same as in fig. 1. The ticks on the y-axis refer to strange matter EoSs with the corresponding $x_{v}$ and B parameters.
The strange matter EoSs were constructed by varying the vector coupling constant ‘$x_{v}$’ and bag parameters ‘B’ by small intervals as discussed in section II.2. Out of the 58 constructed EoSs, the EoS with $x_{v}=0.6$ and B=139.5 possesses the highest evidence value.
Figure 6: Left: EoS plot of the strange matter EoSs; Right: Mass-Radius curves of the strange matter EoSs. The rejected EoSs are plotted in darkblue, while the accepted ones are plotted in salmon.
Figure 7: Mass-Radius curves of the strange matter EoSs. The figure illustrates the rejected and accepted strange matter EoSs when HESS J1731-347 is not considered. The rejected EoSs are plotted in darkblue, while the accepted ones are plotted in salmon.
The fig. 5 shows the odds ratio plot of each quark matter EoS in comparison to the EoS with $x_{v}=0.6$ and B = 139.5. There are 12 EoSs situated in the ‘indecisive’ region, all associated with low values of the bag parameter. The fig. 5 shows a clear trend of increasing evidence against EoSs with increasing values of the bag parameter. We found that 19 EoSs could be decisively rejected, whose bag parameter values were found to be $\geq$ 146.5. Furthermore, for $x_{v}=0.4$, 11 EoSs (with B $\geq$ 144.5) were found to lie in the ’decisive’ region, whereas, for $x_{v}=0.6$, 7 EoSs (with B $\geq$ 146.5) were found to lie in the ‘decisive’ region, implying that by increasing the value of the vector coupling constant, some higher values of the bag parameter could still be preferred by the current observations. The EoSs and their corresponding M-R curves for the rejected and accepted EoS are shown in fig. 6. Similar to the analysis of hadronic EoS, the SM EoSs also prefer stiffer EoS. This again shows that all the current astrophysical observations prefer a stiffer EoS. From fig. 7, we see the difference in MR curves when we do not include HESS J1731-347 data. Without the inclusion of this observation, only a few EoSs were found to be decisively rejected; however, this number significantly increased upon its inclusion.
@@ III.3 Hybrid EoS
The hybrid EoSs constructed (a total of 637 are used in our analysis) in section II.3 were evaluated. XMLSLZ-GM1 with $x_{v}=0.5$ and B = 158.33 was found to have the highest evidence value. In the following sections, we shall denote this EoS as ‘Hyb_best’ to avoid lengthy phrases. The fig. 8 and fig. 14 (left) show the odds ratio plot of each EoS (for different $x_{v}$ values) with respect to Hyb_best.
@@
A total of 429 EoSs were ‘decisively’ rejected when compared with Hyb_best, and there are 6 EoSs with odds ratio values situated in the ‘indecisive’ region. The comparable EoSs are: XMLSLZ-DDME2 ($x_{v}=0.5$, B=158.33), XMLSLZ-DDME2 ($x_{v}=0.44$, B=158.33), XMLSLZ-DDLZ1 ($x_{v}=0.44$, B=158.33), XMLSLZ-MTVTC ($x_{v}=0.44$, B=158.33), RG-SKb ($x_{v}=0.5$, B=155.0) and PT-GRDF2-DD2 ($x_{v}=0.5$, B=158.33). Additionally, it is observed that among the 7 comparable EoSs, including Hyb_best, 6 exhibit a bag parameter value of 158.33, indicating a greater preference for this specific value. However, it should be noted that such a preference is only observed in higher values of the vector coupling constant (fig. 8) and not in lower values. It is also evident in fig. 14 (left) that choosing B=158.33 for $x_{v}=0.2$ does not provide any improvement in the odds ratio value of the EoS family.
In fig. 8 and fig. 14 (left), there are certain EoS families that do not exhibit a lot of change in their evidence values, even after changing the vector coupling constant and bag parameter values. The fig. 9 shows the accepted and rejected hybrid EoSs (left) along with their corresponding M-R curves (right). The discontinuity in FOPT is represented in fig. 11 (left). This reveals that EoSs with early PT are preferred over those with PT at higher pressures.
@@
The fig. 10 illustrates the M-R curves of hybrid EoSs when the data on HESS J1731-347 is not considered. Comparing fig. 9 (right) and fig. 10, we find that inclusion of the information of HESS J1731-347 changes the nature of EoSs which were rejected earlier. Without HESS J1731-347 information, only EoSs having higher radius were favored. Including the data of HESS J1731-347, stars with smaller radii cannot be rejected from our analysis. Analyzing fig. 11 (right) indicates that no comments can be made upon the transition of FOPT without the HESS data, but upon including it (left), one can definitely say that EoSs with early FOPTs are preferred by the observations.
Figure 8: Left: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.44; Right: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.5. Note that in both the plots, EoS refers to Hyb_best, and $EoS_{i}$ refers to the EoS being compared.
Figure 9: Left: EoS plot of the hybrid EoSs; Right: Mass-Radius curves of the hybrid EoSs. The rejected EoSs are plotted in darkblue, while the accepted ones are plotted in salmon.
Figure 10: Mass-Radius curves of the hybrid matter EoSs. The figure illustrates the rejected and accepted hybrid matter EoSs when HESS J1731-347 is not considered. The rejected EoSs are plotted in darkblue, while the accepted ones are plotted in salmon.
Figure 11: Left: The density discontinuity corresponding to the hybrid EoSs from fig. 9 is shown; Right: The density discontinuity corresponding to the hybrid EoSs from fig. 10 is shown. The y-axis represents the pressure, and the x-axis the energy density. The color nomenclature remains the same as previous.
@@ III.4 Comparison among the ‘Indecisive’ EoSs
Figure 12: Odds ratio plot of Hyb_best with the best performing EoSs in each family of EoSs. Note that EoS refers to Hyb_best, and $EoS_{i}$ refers to the EoS being compared.
In the previous subsections, we have found the best performing EoS from the hadronic family, strange matter family, and the hybrid star family EoSs. The odds ratio analysis among the EoSs from each family that lie in the ‘indecisive’ region is performed. The fig. 12 shows the odds ratio plot of each EoS with Hyb_best. It shows that the hybrid EoS Hyb_best performs best among all the EoSs. We find that for the hybrid EoSs, there are no EoSs that lie beyond the ‘indecisive’ region, suggesting that the hybrid EoSs are the most probable among all the families of EoSs. The analysis of the hadronic EoSs reveals that EoSs following the RMF model, the Density-Dependent RMF model, and the Thomas-Fermi approach satisfy the current astrophysical observations the best. For the SM EoSs, we see that the EoSs having a value of B $\leq 139.5$ are most preferred.
@@ IV Summary and Conclusions
The work revisits the current status of the EoSs upon including the mass and radius measurements of the compact object HESS J1731-347 using a bayesian model selection technique. The current astrophysical observations of - PSR J0030+0451, PSR J0740+6620, and GW170817, along with HESS J1731-347 were used to unravel the implications of the latter observation. Starting with constructing three families of EoSs - the first family consists of a hadronic matter comprised mostly of $npe\mu$. The second family of EoSs was built on the modified MIT bag model with vector couplings and comprising three flavored quarks. For the third family of EoSs, we used hybrid EoSs comprising of FOPT.
@@
After obtaining the EoSs, a bayesian model selection was performed for each family of EoSs. In order to evaluate the odds ratio of the EoSs, the evidence integrals were solved by choosing mass as the parameter to be integrated. A uniform prior on mass was adopted, ranging from $0.5M_{\odot}$ to the maximum mass allowed by the EoS. While calculating the odds ratio, we assume each EoS is equally likely, thus reducing the odds ratio to the Bayes factor. The Jeffrey’s scale, which allows to decisively reject EoSs and choose the most suitable EoS models, was used.
@@
The analysis of the hadronic family of EoSs shows that the nuclear EoSs following the Brueckner-Hartree-Fock approximation can be decisively rejected, and while the EoSs based on Effective interactions cannot be effectively rejected; however, model is one of the least effective models in explaining the current astrophysical observations along with the compact object HESS J1731-347. Thus, we conclude that if current observations are composed of hadronic matter, then they are unlikely to have an EoS based on these two models. SPG(M3) (unified NS EoS), following the density-dependent RMF model, was found to be the most likely. The density-dependent RMF model is one of the most effective model that satisfies all the current astrophysical observations best. None of the EoS models coming from the Brussels-Montreal energy density functionals could be decisively rejected. On further investigating the speed of sound of these EoSs, we found that at the intermediate density ranges, a stiffer EoS is preferred by the current astrophysical observations.
@@
Analyzing the SM EoS, it was found that an increase in the Bag value fails to satisfy the current data. EoSs with bag parameter values greater than 146.5 were decisively rejected. Our analysis of the SM EoS shows that if current observations are SSs, then they would inherently prefer an EoS with smaller values of the bag parameter. However, we see that if we increase the value of the vector coupling constant, slightly higher bag values can still be preferred. The EoSs also show a similar trend to the hadronic EoSs, with the stiffer EoSs being preferred over the softer ones.
@@
The analysis of the hybrid EoSs shows that a hybrid EoS referred to as Hyb_best was found to be the most likely hybrid EoS. It was found 6 EoSs comparable to Hyb_best, five of which also have bag parameter values equal to 158.33, indicating a greater preference towards this bag value. However, such a preference is only observed in higher values of the vector coupling constant $x_{v}$. Since the hybrid EoSs comparable to Hyb_best also had a value of B = 158.33, it suggests that contrary to our SM EoSs, including HESS J1731-347 in our analysis shows a preference for FOPT with larger Bag values. It was also found that the EoSs with FOPT at higher densities were likely disfavored compared to EoSs having an early PT.
Further comparison among all the families of EoSs that lie in the ‘indecisive’ region of our analysis, we find that all the hybrid EoSs perform the best, suggesting that explaining current astrophysical observations using a hybrid EoS is the most likely scenario. The SM EoSs also hint that the value of the bag parameter should be less than 139.5, such that all the astrophysical observations are satisfied best. We could not decisively comment on the hadronic EoS family. However, we found that some EoSs based on the RMF model, Density-Dependent RMF model, and Thomas-Fermi approximation are comparable to Hyb_best.
@@
It is to be noted that since we have considered only the modified MIT Bag model with vector interactions as the quark counterpart in our hybrid EoSs, hence there also lies a possibility that the rejected hybrid EoSs might be able to explain the current observations upon considering different types of quark counterparts.
@@ Acknowledgement
The authors would also like to thank IISER Bhopal for providing the infrastructure for this work. The authors also acknowledge Prasanta Char for their helpful discussions. DK acknowledges Bhaskar Biswas for email correspondence. ST would like to thank Akshat Singh, Pratik Thakur, Shamim Haque, and Rishi Gupta for their helpful discussions. SC acknowledges the Prime Minister’s Research Fellowship (PMRF), Ministry of Education Govt. of India, for a graduate fellowship. RM and DK acknowledge the Science and Engineering Research Board (SERB), Govt. of India, for monetary support in the form of a Core Research Grant (CRG/2022/000663).
@@
The data can be availed upon reasonable request to the authors.
appendix
IV.1 Status of EoSs without HESS J1731-347
Figure 13: Left: Odds ratio plot of the hadronic matter EoS PCP(BSK25) with other hadronic matter EoSs. These are the results obtained when HESS J1731-347 was not taken into consideration. The color scheme is the same as in fig. 1; Right: Odds ratio plot of the strange matter EoS with $x_{v}=0.4$ and B$=139.0$ with other strange matter EoSs. These are the results obtained when HESS J1731-347 was not taken into consideration. The color scheme is the same as in fig. 1. The ticks on the y-axis refer to strange matter EoSs with the corresponding xv and B parameters.
In this section, we discuss the status of the EoSs in the context of astrophysical observations of only: PSR J0740+6620, PSR J0030+0451, and GW170817. Only including these observations, we illustrate how each family of hadronic, quark, and hybrid family of EoSs behave in fig. 13, fig. 14 (right), and fig. 15. In fig. 13 we show the odds ratio plot for the hadronic and SM EoSs. From the hadronic family of EoSs, we observe that none of the EoSs could be decisively rejected contrary to the results discussed upon including HESS observation. Furthermore, in this scenario, PCP(BSK25) (following the Brussels-Montreal energy density functional) was found to be the most likely EoS, with 29 other EoSs lying in the ’indecisive’ region. This is a significantly greater number than when the data of HESS is taken into consideration.
The SM EoS analysis shows that though there are several EoSs that can be decisively rejected based only on the other observations, however, the total number of EoSs that can be decisively rejected upon including HESS data significantly increases. In this scenario, the EoS with $x_{v}$ = 0.4 and B = 139.0 is found to be the most likely, with 17 EoSs lying in the ’indecisive’ region. Therefore, a similar trend is also followed by comparable equations, where the number of comparable equations decreases when the data of HESS J1731-347 is included.
Figure 14: Left: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.2 and when HESS was considered; Right: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.2 and HESS was not considered. Note that in both the plots, EoS refers to Hyb_best, and $EoS_{i}$ refers to the EoS being compared.
@@ IV.1 Status of EoSs without HESS J1731-347
Figure 15: The figure illustrates the odds ratio plot obtained when HESS J1731-347 was not considered. Left: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.44; Right: Odds ratio plot of Hyb_best with EoSs having fixed $x_{v}$ value of 0.5. Note that in both the plots, EoS refers to Hyb_best, and $EoS_{i}$ refers to the EoS being compared.

### v2

@@ III.2 Strange EoS
Figure 3: Odds ratio plot of the strange matter EoS with $x_{v}=0.4$ and B = 139.0 with other strange matter EoSs. The colour scheme is the same as in fig. 1. The ticks on the y-axis refer to strange matter EoSs with the corresponding $x_{v}$ and B parameters. $EoS_{i}$ refers to the equation of state being compared.
The strange matter EoSs were constructed by varying the scaled coupling constant ‘$x_{v}$’ and bag parameters ‘B’ by small intervals as discussed in section II.2. Out of the 58 constructed EoSs, the EoS with $x_{v}=0.4$ and B=139.0 possesses the highest evidence value.
Figure 4: (a): EoS plot of the strange matter EoSs; (b): Mass-Radius curves of the strange matter EoSs; (c): Mass-Radius curves of the strange matter EoSs when HESS J1731-347 is not considered. The colour nomenclature remains the same as the previous one.
Fig. 3 shows the odds ratio plot of each quark matter EoS in comparison to the EoS with $x_{v}=0.4$ and B = 139.0. There are 12 other EoSs situated in the ‘indecisive’ region, all associated with low values of the bag parameter. Fig. 3 shows a clear trend of increasing evidence against EoSs with increasing values of the bag parameter, with only two minor deviations from it. This should be attributed to the fact that increasing the value of the bag parameter decreases the stiffness of the EoS. Therefore, our analysis suggests that similar to the hadronic matter EoSs, strange matter EoSs also prefer stiff EoSs.
We found that 17 EoSs could be decisively rejected, whose bag parameter values were found to be $\geq$ 146.5. Furthermore, for $x_{v}=0.4$, 10 EoSs (with B $\geq$ 145.0) were found to lie in the ’decisive’ region, whereas, for $x_{v}=0.6$, 7 EoSs (with B $\geq$ 146.5) were found to lie in the ‘decisive’ region, implying that by increasing the value of the scaled coupling constant, some higher values of the bag parameter could still be preferred by the current observations. This is an expected result since an increase in the value of the scaled coupling constant increases the stiffness of the EoS.
The rejected and accepted EoSs and their corresponding M-R curves are shown in figs. 4(a) and 4(b) respectively. In the M-R plot, the EoSs that barely cut the PSR J0030+0451 contour are rejected, and those that have a significant overlap with the contour are accepted by our analysis. Looking at fig. 4(a), it can be seen that at low pressure, there is a cutoff value of density for the EoS to be rejected (approximately 225 MeVfm-3).
From figs. 4(b) and 4(c), we see the difference in MR curves when we do not include HESS J1731-347 data. Without the inclusion of this observation, only a few EoSs were found to be decisively rejected; however, this number significantly increased upon its inclusion.
@@ III.3 Hybrid EoS
The hybrid EoSs constructed (545 are used in our analysis) in section II.3 were evaluated. XMLSLZ-DDME2 with $x_{v}=0.44$ and B = 158.33 was found to have the highest evidence value. In the following sections, we shall denote this EoS as ‘Hyb_best’ to avoid lengthy phrases. Figs. 5 and LABEL:fig:Hybrid_xv02 show the odds ratio plot of each EoS (for different $x_{v}$ values) with respect to Hyb_best.
@@
A total of 351 EoSs were ‘decisively’ rejected when compared with Hyb_best, and there are 4 other EoSs with odds ratio values situated in the ‘indecisive’ region. The comparable EoSs are: XMLSLZ-DDME2 ($x_{v}=0.5$, B=158.33), XMLSLZ-DDLZ1 ($x_{v}=0.44$, B=158.33), RG-SKb ($x_{v}=0.5$, B=155.0) and PT-GRDF2-DD2 ($x_{v}=0.5$, B=158.33). Additionally, it is observed that among the five comparable EoSs, including Hyb_best, four of them exhibit a bag parameter value of 158.33, indicating a greater preference for this specific value. However, it should be noted that such a preference is only observed in higher values of the scaled coupling constant (fig. 5) and not in lower values. It is also evident in LABEL:fig:Hybrid_xv02 that choosing B=158.33 for $x_{v}=0.2$ does not improve the odds ratio value of the EoS family.
In figs. 5 and LABEL:fig:Hybrid_xv02, certain EoS families do not exhibit much change in their evidence values, even after changing the scaled coupling constant and bag parameter values. Figs. 7(a) and 7(b) show the accepted and rejected hybrid EoSs and their corresponding M-R curves, respectively. Since it is difficult to comment on the hybrid EoSs using only fig. 7(b), we utilize the two parameters of Maxwell’s construction: width of the discontinuity and the onset of phase transition.
Fig. 8(a) illustrates both these parameters for the rejected and accepted EoSs. It can be seen that EoSs with an early onset of PT are preferred over those with a PT at higher densities. We conclude that the onset density of PT is more important, and the preference of the EoS does not depend on the width of the discontinuity.
@@
Fig. 7(c) illustrates the M-R curves of hybrid EoSs when the data on HESS J1731-347 is not considered. Comparing fig. 7(b) and fig. 7(c), we find that the inclusion of the information of HESS J1731-347 dramatically increases the number of EoSs being rejected and puts a better constraint on the nature of the EoS. Without HESS J1731-347 information, only EoSs having higher radii were favoured. Including the data of HESS J1731-347, stars with smaller radii cannot be rejected from our analysis. Analyzing fig. 8(b) indicates that no comments can be made upon the transition of FOPT without the HESS data. However, upon including it (fig. 8(a)), one can definitely say that EoSs with early FOPTs are preferred by the observations.
@@ III.4 Comparison among the ‘Indecisive’ EoSs
In the previous subsections, we have found the best performing EoS from the hadronic family, strange matter family, and the hybrid star family of EoSs. In this section, the odds ratio analysis among the EoSs from each family that lie in the ‘indecisive’ region is performed. Fig. 9 shows the odds ratio plot of each EoS with Hyb_best. It shows that the hybrid EoS Hyb_best performs best among all the EoSs. We find that for the hybrid EoSs, no EoSs lie beyond the ‘indecisive’ region, suggesting that the hybrid EoSs are the most probable among all the families of EoSs. The analysis of the hadronic EoSs reveals that EoSs following the RMF model, the Density-Dependent RMF model, and the Thomas-Fermi approach satisfy the current astrophysical observations the best. For the SM EoSs, we see that the EoSs having a value of B $\leq 139.5$ are most preferred.
Figure 5: (a): Odds ratio plot of Hyb_best with EoSs having a fixed $x_{v}$ value of 0.44; (b): Odds ratio plot of Hyb_best with EoSs having a fixed $x_{v}$ value of 0.5. The colour nomenclature of the disks is the same as in figs. 1 and 3. Each sector belongs to a hadronic family. The different shaped points correspond to different bag values referenced in (a). The different colours of the points help distinguish continuous sets of hadronic families. The odds ratio value of each EoS acts as the radius value in the plot. This value is then plotted in the sector of the hadronic family of that EoS, with the shape of its corresponding bag value.
Figure 7: (a): EoS plot of the hybrid EoSs; (b): Mass-Radius curves of the hybrid EoSs; (c): Mass-Radius curves of the hybrid EoSs when HESS J1731-347 is excluded. The colour nomenclature remains the same as the previous one.
Figure 8: (a): The density discontinuity corresponding to the hybrid EoSs from fig. 7(b) is shown; (b): The density discontinuity corresponding to the hybrid EoSs from fig. 7(c) is shown. The colour nomenclature remains the same as the previous one.
Figure 9: Odds ratio plot of Hyb_best with the best performing EoSs in each family of EoSs. Note that EoS above refers to Hyb_best, and $EoS_{i}$ refers to the equation of state being compared.
@@ IV Summary and Conclusions
The work improves the current status of the EoSs upon including the mass and radius measurements of the compact object HESS J1731-347 using a Bayesian model selection technique. The current astrophysical observations of - PSR J0030+0451, PSR J0740+6620, J0437-4715, and GW170817, along with HESS J1731-347, were used to unravel the implications of the latter observation. Starting with constructing three families of EoSs - the first family consists of hadronic matter. The second family of EoSs was built on the modified MIT bag model with scaled couplings and comprising of three flavoured quarks. For the third family of EoSs, we used hybrid EoSs comprising a FOPT.
@@
After obtaining the EoSs, a Bayesian model selection was performed for each family of EoSs. In order to evaluate the odds ratio of the EoSs, the evidence integrals were solved by choosing mass as the parameter to be integrated. A uniform prior on mass was adopted, ranging from $0.5M_{\odot}$ to the maximum mass allowed by the EoS. While calculating the odds ratio, we assume each EoS is equally likely, thus reducing the odds ratio to the Bayes factor. We adopted Jeffrey’s scale, allowing us to decisively reject EoSs and choose the most suitable EoS models.
@@
The analysis of the hadronic family of EoSs shows that the nuclear EoSs following the Brueckner-Hartree-Fock approximation and EoSs based on effective interactions are least effective in explaining the current astrophysical observations along with the compact object HESS J1731-347. XMLSLZ(DD-LZ1), following the Thomas Fermi approximation, was found to be the best performing EoS. Density-dependent RMF models and the Thomas Fermi approximation are the most effective models in explaining the current observations.
Most of the EoSs were accepted or rejected based on either the radius or the tidal deformability bound of the observations. The accepted EoSs showed a back-bending effect on the MR curve, which is due to a sudden stiffening of the EoS. Moreover, these models showed a non-monotonous nature in their speed of sound, with few showing maxima, which can be associated with the emergence of new degrees of freedom at higher densities.
@@
Analyzing the SM EoS, it was found that an increase in the Bag value fails to satisfy the current data as it effectively makes the EoS soft. Our analysis of the SM EoS shows that if current observations are SSs, they would inherently prefer an EoS with smaller values of the bag parameter. However, if we increase the value of the scaled coupling constant, slightly higher bag values can still be preferred. There is a clear distinction between the rejected and accepted SM EoSs, both in the EoS and MR plot, with the data of the pulsar PSR J0030+0451 playing a decisive role in the segregation of the SM EoSs. Additionally, the analysis hints towards an energy density cut-off value at low pressures. If the energy density of an EoS at low pressure lies above this threshold value, it is rejected by our analysis.
@@
The analysis of the hybrid EoSs shows that a hybrid EoS referred to as Hyb_best is the most likely hybrid EoS having a bag parameter value equal to 158.33 and scaled coupling constant value equal to 0.44. It was found that 4 other EoSs are comparable to Hyb_best, three of which also have bag parameter values equal to 158.33, indicating a greater preference towards this bag value. However, such a preference is only observed in higher values of the scaled coupling constant $x_{v}$. Although the MR curve does help in distinguishing the accepted EoSs from the rejected EoSs, this segregation is more dependent on the onset density of phase transition, preferring early PT densities.
On comparing all the EoSs of different families that lie in the ‘indecisive’ region of our analysis, we find that all the hybrid EoSs perform the best, suggesting that explaining current astrophysical observations using a hybrid EoS is the most likely scenario.
The results are consistent with the recent observation that EoSs show a non-monotonic speed of sound. The non-monotonicity is usually associated with the appearance of extra degrees of freedom (in this case, quark matter), suggesting a phase transition at intermediate densities in neutron stars. This is reflected in Hyb_best being the most favored EoS.
@@
It is to be noted that we have considered only the modified MIT Bag model with vector interactions as the quark counterpart in our hybrid EoSs, hence there lies a possibility that the rejected hybrid EoSs might be able to explain the current observations upon considering different types of quark counterparts. Since the current analysis does not consider the Gibbs construction, there also lies a possibility of explaining the present astrophysical observations using a FOPT constructed using the Gibbs construction. All such extensions are our future endeavors.
@@ Acknowledgement
The authors would also like to thank IISER Bhopal for providing the infrastructure for this work. The authors also acknowledge Prasanta Char for their helpful discussions. DK acknowledges Bhaskar Biswas for email correspondence. ST would like to thank Akshat Singh, Pratik Thakur, Shamim Haque, and Rishi Gupta for their helpful discussions. SC acknowledges the Prime Minister’s Research Fellowship (PMRF), Ministry of Education Govt. of India, for a graduate fellowship. RM and DK acknowledge the Science and Engineering Research Board (SERB), Govt. of India, for monetary support through a Core Research Grant (CRG/2022/000663).
@@
The data can be made available upon reasonable request by the authors.
Appendix A:   Effect of the CET Constraint
In section II, we discussed the construction of the CET band in order to use it to filter EoSs. This section discusses the changes in our analysis when the CET constraint is not considered. Initially, our analysis included 50 hadronic EoSs and 637 hybrid EoSs. We found that from the 50 hadronic EoSs considered, 12 of them did not satisfy the CET bound.
On performing the entire analysis with these EoSs, we saw that 8 out of the 12 EoSs that did not satisfy the CET constraint were already rejected from our analysis. Of the remaining 4 EoSs, two were in the ‘substantial’ regime, and two were in the ‘strong’ regime. Similarly, on performing the analysis for the initial 637 hybrid EoSs, we found that 93 EoSs did not satisfy the CET constraint. Out of these 92 EoSs, 86 were rejected from our analysis, 2 were in the ‘strong’ regime, 3 were in the ‘substantial’ regime, and 1 was found to be in the ‘indecisive’ regime.
Appendix B:   Status of EoSs without HESS J1731-347
Figure 10: (a): Odds ratio plot of the hadronic matter EoS PT(GRDF2-DD2) with other hadronic matter EoSs. These are the results obtained when HESS J1731-347 was not taken into consideration. The colour scheme is the same as in fig. 1; (b): Odds ratio plot of the strange matter EoS with $x_{v}=0.4$ and B$=139.0$ with other strange matter EoSs. These are the results obtained when HESS J1731-347 was not taken into consideration. The colour scheme is the same as in fig. 1. The ticks on the y-axis refer to strange matter EoSs with the corresponding xv and B parameters.
In this section, we discuss the status of the EoSs in the context of astrophysical observations of only PSR J0740+6620, PSR J0030+0451, PSR J0437-4715, and GW170817. Only including these observations, we illustrate how each of the hadronic, quark, and hybrid family of EoSs behave in figs. 10, LABEL:fig:Hybrid_pulsarplot and LABEL:fig:Hybrid_xv02_onlypulsar. In fig. 10, we show the odds ratio plot for the hadronic and SM EoSs. From the hadronic family of EoSs, we observe that none of the EoSs could be decisively rejected contrary to the results discussed upon including HESS observation. Furthermore, in this scenario, PT(GRDF2-DD2) (following the Brussels-Montreal energy density functional) was the most likely EoS, with 24 other EoSs lying in the ’indecisive’ region. This is a significantly greater number than when the data of HESS is considered.
The SM EoS analysis shows that although several EoSs can be decisively rejected based only on these observations, the total number of EoSs that can be decisively rejected upon including HESS data significantly increases. In this scenario, the EoS with $x_{v}$ = 0.4 and B = 139.0 is the most likely, with 15 other EoSs lying in the ’indecisive’ region. Therefore, a similar trend is also followed by comparable equations, where the number of comparable equations decreases when the data of HESS J1731-347 is included.